# photometric stereo; multi-scale features; deep neural networks; attention mechanisms

### v1

@@ Abstract
keywords:
photometric stereo; multi-scale features; deep neural networks; attention mechanisms
[…]
@@ 1 Introduction
Since the creation of the first photometric stereo (PS) algorithm by Woodham Woodham (1980) under the Lambert hypothesis, acquiring images with varying light directions using linear response cameras and utilizing the PS algorithm to obtain accurate normal maps of objects have been a focus of researchers, especially in areas where the object’s structure and texture have undergone changes Sun and Lam (2011); Nie et al. (2020); Jian et al. (2019). Compared to traditional methods, deep neural networks have the capability to imitate intricate global lighting effects that cannot be represented by previous mathematical formulas, resulting in significantly enhanced accuracy of the results. However, unlike other computer vision tasks that often have a fixed input size or sequence, photometric stereo networks require handling an unknown order problem, making it difficult for Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) to handle due to their limited flexibility.
@@
To address the order-agnostic nature of the photometric stereo task and produce an accurate normal map in a complex region, previous studies have proposed several approaches. These include using fused feature maps from max-pooling operations or fixed-size observation maps as inputs to neural networks, increasing the network’s depth to improve feature representation, and generating virtual training datasets with varying complexity levels to enhance the network’s fitting ability Chen et al. (2018); Ikehata (2018); Logothetis et al. (2021). In summary, existing approaches primarily focus on providing various solutions to intricate global illumination issues, including but not limited to shadows and specular highlights. While these approaches have improved the accuracy of normal map prediction, they still face difficulties in dealing with “difficult” regions, such as the structure changing rapidly and the material changes, due to the lack of integration of different stages’ features of the input image.
@@
Recently, research has shown that low-level vision tasks, including photometric stereo, occur at multiple scales in natural scenes and require careful consideration of scale information Ju et al. (2022); Liu et al. (2022). On the contrary, many existing methods rely on complex network structures and ignore the fact that scale information changes the focus of one’s observation of an image. For example, high-resolution images input at the initial stage of the network can extract rich texture details that help to focus on complex structural regions, whereas low-resolution images at deeper levels tend to emphasize contour information, which facilitates the identification of surfaces with spatially varying materials. Therefore, extracting multi-sale image features from different stages is crucial to solving the problem of fuzzy details in normal graphs.
@@
In addition to extracting multi-scale features, since the high-dimensional feature map of the image contains many characteristics of the object surface, attention mechanisms de Santana Correia and Colombini (2022); Woo et al. (2018) can also be used to adaptively adjust feature weights while preserving all extracted features, enabling the network to focus on specific local areas such as material changes or areas affected by global light for optimized feature extraction. Such attention mechanisms have been shown to improve the expressive power of the network and enable it to treat different areas of the image differently, thereby improving the overall performance of the photometric stereo network. Combined with this theory, spatial attention assigns higher weight parameters to areas that require greater attention, such as surface structure or material changes, while the channel attention enhances the weight of channels related to normal regression and suppresses irrelevant channels, for example, light intensity or roughness-related channels, allowing the network to retain the most useful information with respect to predicting the normal vector of the “difficult” region.
@@
This paper proposes a novel photometric stereo network, the residual multi-scale attention feature fusion photometric stereo network (RMAFF-PSN), which effectively integrates multi-scale feature information from both shallow and deep stages using attention-weighted fusion. As shown in the boxes in Figure 1, the RMAFF-PSN can effectively deal with the challenges posed by both retaining salient details of object material variations and structurally complex regions during network propagation. To achieve this, we constructed a residual multi-scale attention feature fusion module inspired by previous work on residual multi-scale network structure He et al. (2016); Gao et al. (2019); Zhu et al. (2022). To process the input image, we extract feature information separately from its high- and low-resolution versions and then stitch together the shallow and deep features. The residual multi-scale attention feature fusion module leverages multiple view fields across various scales to enhance multi-scale feature information, in a similar manner to residual connections. Subsequently, the attention mechanism optimizes the feature maps along the channel and spatial dimensions, followed by feature combination with the shallow-deep feature information to complete the steps of feature extraction, enhancement, and optimization. This approach enables the network to effectively capture characteristic information that best describes the different areas of an object’s surface, so as to reconstruct the rich and accurate surface normal map.
@@ 2 Related Work
Deep learning has demonstrated remarkable efficacy in the field of optical image processing. Santo Santo et al. (2017) was the first to apply deep learning methods to photometric stereo, achieving prediction accuracies that significantly surpassed those of traditional algorithms. Currently, the prevailing view is that deep learning approaches are predominantly used to tackle photometric stereo tasks using per-pixel and all-pixel methods Zheng et al. (2020); Ju et al. (2022).
@@
Both the per-pixel and all-pixel methods have their respective drawbacks. The former ignores internal image information and is limited by the size setting of the observation map, while the latter overuses 3 $\times$ 3 convolutions and may miss important details. In response to these issues, Yao Yao et al. (2020) proposes using SGC filters to extract feature information from topologically adjacent points, and combines these two methods in a sequential manner. Ju Ju et al. (2021) adopts a self-supervised approach that learns the attention-weighted loss for each pixel point and introduces penalties for different surface areas, which can retain more detailed gradient information.
@@
Here $I_{j}$ represents the image pixel intensity in the $jth$ illumination direction. $\rho$ is the bidirectional reflectance distribution function (BRDF, an important formula in the field of optics and graphics used to describe how light is reflected from a given direction of incident light and outgoing light at a surface Jensen et al. (2001)) and $max(\bm{n}^{T}\bm{l_{j}},0)$ represents an attached shadow, and $\mu_{j}$ represents the noise of the camera and the environment. We assume that the pixel intensities of an image are normalized by the corresponding lighting intensities, so $\bm{l}$ can be viewed as a unit vector. The challenge of dealing with non-Lambertian surfaces limits the applicability of some photometric stereo methods that are suitable only for Lambertian surfaces, such as the least squares method. To address this issue, deep learning methods have emerged as a powerful tool for fitting surfaces of objects that cannot be expressed by mathematical formulas. However, even with the impressive fitting ability of deep learning methods, reconstruction results can still be fuzzy and have large angle errors due to the shape or material of some areas. As shown in Figure 2, complex scenes with intricate structures can be particularly challenging. Fortunately, this paper proposes a novel approach to address these “difficult” regions using multi-scale attention feature fusion, which is both simple and effective.
@@ 3 Methods
In the normal regression part, to prevent overfitting caused by a deep network, we added a dense-block structure Huang et al. (2017). This structure enhances feature propagation and reuses features from low-dimensional inputs to improve the accuracy of the network in predicting the normal image pixels while retaining the features of shallow local regions. In addition, we used batch normalization between each layer of the convolutional neural network to adjust the weights of the neurons into a standard normal distribution regularization layer. Our network architecture consists of six convolutional layers, two downsampling layers, five upsampling layers, two residual multi-scale attention feature fusion modules, one max-pooling layer, and one dense-block module. Furthermore, we utilized L2-Norm layers to normalize the surface normal vectors.
@@
$\bm{Branch_{j}^{i}}=\left\{\begin{array}[]{l}AsyConv(\bm{F_{i}}){\rm{}},{\rm{}}j=1\\ AsyConv(\bm{F_{i}}\oplus\bm{Branc{h_{j-1}}}){\rm{}},{\rm{}}j=2,3,4\end{array}\right.$ (3)
@@ 4 Experimental Results
4.1 Datasets
The insufficient number of datasets is a limiting factor that affects the accuracy of photometric stereo predictions. In the experiments conducted in this study, the training datasets used were the same as those used by most all-pixel methods. The two shape datasets are the Blobby and Sculpture datasets Johnson and Adelson (2011); Wiles and Zisserman (2017), which were rendered using the MERL BRDF dataset Matusik (2003) containing information on 100 different materials. The two datasets contain 25,920 and 59,292 shapes, respectively, and each captured shape has 64 different lighting conditions, resulting in a total of 5.4 million images. The Sculpture dataset is more complex and contains more detailed information. During the training process, we used a ratio of 99:1 for training and validation data.
The DiLiGenT dataset Shi et al. (2016) consists of 10 real-world objects with complex shapes and different materials, each captured in 96 lighting orientations. Ground truth information is provided for each object in the dataset. To evaluate the effectiveness of our proposed model, we used this dataset for quantitative assessment.
@@ 4.2 Network Analysis
4.2.1 Effectiveness of a Residual Multi-Scale Attention Feature Fusion Module
@@
“w/o RMAFF” removes the RMAFF module from the network and only concatenates shallow-deep features. “Single RMAFF” employs a single branch RMFE module to extract different levels of features from the deep image. “w/o Attention” removes the attention allocation mechanism from the module and directly connects the feature information extracted by the residual multi-scale module. “RMAFF+AFF” replaces the concat operation with the AFF structure Dai et al. (2021) for feature stitching. Our approach uses two RMFE modules to extract different features from the shallow and deep layers and then stitches them together using the concat operation.
@@
4.2.2 The Validity of Normal Regression Network Structures
@@
4.2.3 Effect of Different Resolutions and Number of Input Images
Images of different resolutions may contain varying levels of information, with larger images often containing richer features such as geometry and texture Sun et al. (2022). Our proposed RMAFF module is designed to fully perceive light and dark variations within an image and extract the most appropriate image features based on those variations. However, due to GPU memory limitations, training models with larger numbers and higher resolutions of input images can significantly increase the training time.
@@ 4.2.3 Effect of Different Resolutions and Number of Input Images
4.2.4 Effect of Different Training Datasets
In this study, the influence of dataset complexity on the accuracy of the proposed network model is analyzed. It is well known that the accuracy of the network model can be affected by the complexity of the dataset, as even better networks may not necessarily improve model accuracy, a phenomenon known as “Kolmogorov complexity” in machine learning Bolon-Canedo and Remeseiro (2020); Kabir and Garg (2023). To investigate the effect of dataset complexity on network performance, we conducted separate experiments on three different training datasets and reconstructed the normal map of the “Buddha” object. It should be noted that everything was kept the same for the network except for the training dataset. The experimental results are shown in Figure 9. Our analysis indicates that the network’s performance in reconstructing surface normals is better when the surface complexity of the training dataset is higher (Train_Blobby vs. Train_Sculpture) and the dataset has a larger number of samples (Train_Blobby vs. Train_Blobby+Sculpture). Therefore, we conclude that the complexity of the object shape and material is essential in the photometric stereo task, enabling the network to capture more features of the object surface and achieve a more accurate normal estimation during the test phase.
@@ 4.2.4 Effect of Different Training Datasets
4.3 Benchmark Comparison of the DiLiGenT Dataset
We conducted a comprehensive comparison of the RMAFF-PSN with several commonly used methods, which included linear least squares-based methods (L2 Woodham (1980)), four per-pixel methods (CNN-PS Ikehata (2018), LMPS Li et al. (2019), SPLINE-Net Zheng et al. (2019), PX-Net Logothetis et al. (2021)), and four all-pixel methods (IRPS Taniai and Maehara (2018), PS-FCN Chen et al. (2018), CHR-PSN Ju et al. (2022), MF-PSN Liu et al. (2022)). To evaluate the performance of each method, we used the DiLiGenT dataset, which included 96 input images for all methods, except for the “Bear” object, which used 76 images due to corruption in the first 20 images. The results of our experiments are presented in Table 2. Our proposed RMAFF-PSN method outperformed most of the networks, exhibiting higher accuracy than the existing deep learning methods under varying light distributions and image sizes (Figure 8b,d). Furthermore, we provide a visual comparison of the predicted normal maps for each object in the DiLiGenT dataset in Figure 10.
@@ 4.3 Benchmark Comparison of the DiLiGenT Dataset
L2 Woodham (1980) 4.10 8.41 8.89 8.39 14.65 14.92 18.50 19.80 25.60 30.62 15.39
WG10 Wu et al. (2011) 2.06 6.73 7.18 6.50 13.12 10.91 15.70 15.39 25.89 30.01 13.35
AZ08 Alldrin et al. (2008) 2.71 6.53 7.23 5.96 11.03 12.54 13.93 14.17 21.48 30.50 12.61
GC10 Goldman et al. (2009) 3.21 8.22 8.53 6.62 7.90 14.85 14.22 19.07 9.55 27.84 12.00
IA14 Ikehata and Aizawa (2014) 3.34 6.74 6.64 7.11 8.77 10.47 9.71 14.19 13.05 25.95 10.60
ST14 Shi et al. (2013) 1.74 6.12 6.51 6.12 8.78 10.60 10.09 13.63 13.93 25.44 10.30
SPLINE-Net Zheng et al. (2019) 4.51 6.49 8.29 5.28 10.89 10.36 9.62 15.50 7.44 17.93 9.63
EW20 Enomoto et al. (2020) 1.58 6.30 6.67 6.38 7.26 13.69 11.42 15.49 7.80 18.74 9.53
DPSN Wiles and Zisserman (2017) 2.02 6.54 7.05 6.31 7.86 12.68 11.28 15.51 8.01 16.86 9.41
CK18 Simchony et al. (1990) 1.50 5.74 6.24 4.97 8.64 8.86 10.00 11.44 11.33 21.90 9.06
IRPS Taniai and Maehara (2018) 1.47 5.44 6.09 5.79 7.76 10.36 11.47 11.03 6.32 22.59 8.83
LMPS Li et al. (2019) 2.40 6.11 6.54 5.23 7.48 9.89 8.61 13.68 7.98 16.18 8.41
PS-FCN Chen et al. (2018) 2.82 6.16 7.13 7.55 7.25 7.91 8.60 13.33 7.33 15.85 8.39
Attention-PSN Ren et al. (2022) 2.93 6.14 6.92 4.86 6.97 7.75 8.42 12.90 6.86 15.44 7.92
DR-PSN Ju et al. (2021) 2.27 5.42 7.08 5.46 7.21 7.84 8.49 12.74 7.01 15.40 7.90
GPS-Net Alldrin et al. (2008) 2.92 5.42 6.04 5.07 7.01 7.77 9.00 13.58 6.14 15.14 7.81
CHR-PSN Ju et al. (2022) 2.26 5.97 7.04 6.35 6.76 7.15 8.32 12.52 6.05 15.32 7.77
CNN-PS Ikehata (2018) 2.12 4.38 5.37 4.20 6.38 8.07 7.42 12.12 7.92 13.83 7.18
PS-FCN(Norm.) Chen et al. (2020) 2.67 4.76 6.17 7.72 7.15 7.53 7.84 10.92 6.72 12.39 7.39
PX-Net Logothetis et al. (2021) 2.03 4.39 5.08 4.13 5.10 7.61 6.90 10.26 4.69 13.10 6.33
NormAttention-PSN Ju et al. (2022) 2.93 4.65 5.96 5.48 6.42 7.12 7.49 9.93 5.99 12.28 6.83
@@
4.4 Qualitative Comparison of Other Real-World Datasets

### v2

# RMAFF-PSN: A Residual Multi-Scale Attention Feature Fusion Photometric Stereo Network

@@ 1 Introduction
Since the creation of the first photometric stereo (PS) algorithm by Woodham [1] under the Lambert hypothesis, acquiring images with varying light directions using linear response cameras and utilizing the PS algorithm to obtain accurate normal maps of objects have been a focus of researchers, especially in areas where the object’s structure and texture have undergone changes [2, 3, 4]. Compared to traditional methods, deep neural networks have the capability to imitate intricate global lighting effects that cannot be represented by previous mathematical formulas, resulting in significantly enhanced accuracy of the results. However, unlike other computer vision tasks that often have a fixed input size or sequence, photometric stereo networks require handling an unknown order problem, making it difficult for Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) to handle due to their limited flexibility.
@@
To address the order-agnostic nature of the photometric stereo task and produce an accurate normal map in a complex region, previous studies have proposed several approaches. These include using fused feature maps from max-pooling operations or fixed-size observation maps as inputs to neural networks, increasing the network’s depth to improve feature representation, and generating virtual training datasets with varying complexity levels to enhance the network’s fitting ability [5, 6, 7]. In summary, existing approaches primarily focus on providing various solutions to intricate global illumination issues, including but not limited to shadows and specular highlights. While these approaches have improved the accuracy of normal map prediction, they still face difficulties in dealing with “difficult” regions, such as the structure changing rapidly and the material changes, due to the lack of integration of different stages’ features of the input image.
@@
Recently, research has shown that low-level vision tasks, including photometric stereo, occur at multiple scales in natural scenes and require careful consideration of scale information [8, 9]. On the contrary, many existing methods rely on complex network structures and ignore the fact that scale information changes the focus of one’s observation of an image. For example, high-resolution images input at the initial stage of the network can extract rich texture details that help to focus on complex structural regions, whereas low-resolution images at deeper levels tend to emphasize contour information, which facilitates the identification of surfaces with spatially varying materials. Therefore, extracting multi-sale image features from different stages is crucial to solving the problem of fuzzy details in normal graphs.
@@
In addition to extracting multi-scale features, since the high-dimensional feature map of the image contains many characteristics of the object surface, attention mechanisms [10, 11] can also be used to adaptively adjust feature weights while preserving all extracted features, enabling the network to focus on specific local areas such as material changes or areas affected by global light for optimized feature extraction. Such attention mechanisms have been shown to improve the expressive power of the network and enable it to treat different areas of the image differently, thereby improving the overall performance of the photometric stereo network. Combined with this theory, spatial attention assigns higher weight parameters to areas that require greater attention, such as surface structure or material changes, while the channel attention enhances the weight of channels related to normal regression and suppresses irrelevant channels, for example, light intensity or roughness-related channels, allowing the network to retain the most useful information with respect to predicting the normal vector of the “difficult” region.
@@
This paper proposes a novel photometric stereo network, the residual multi-scale attention feature fusion photometric stereo network (RMAFF-PSN), which effectively integrates multi-scale feature information from both shallow and deep stages using attention-weighted fusion. As shown in the boxes in Figure 1, the RMAFF-PSN can effectively deal with the challenges posed by both retaining salient details of object material variations and structurally complex regions during network propagation. To achieve this, we constructed a residual multi-scale attention feature fusion module inspired by previous work on residual multi-scale network structure [12, 13, 14]. To process the input image, we extract feature information separately from its high- and low-resolution versions and then stitch together the shallow and deep features. The residual multi-scale attention feature fusion module leverages multiple view fields across various scales to enhance multi-scale feature information, in a similar manner to residual connections. Subsequently, the attention mechanism optimizes the feature maps along the channel and spatial dimensions, followed by feature combination with the shallow-deep feature information to complete the steps of feature extraction, enhancement, and optimization. This approach enables the network to effectively capture characteristic information that best describes the different areas of an object’s surface, so as to reconstruct the rich and accurate surface normal map.
@@ 2 Related Work
Deep learning has demonstrated remarkable efficacy in the field of optical image processing. Santo [15] was the first to apply deep learning methods to photometric stereo, achieving prediction accuracies that significantly surpassed those of traditional algorithms. Currently, the prevailing view is that deep learning approaches are predominantly used to tackle photometric stereo tasks using per-pixel and all-pixel methods [16, 17].
@@
Both the per-pixel and all-pixel methods have their respective drawbacks. The former ignores internal image information and is limited by the size setting of the observation map, while the latter overuses 3 $\times$ 3 convolutions and may miss important details. In response to these issues, Yao [18] proposes using SGC filters to extract feature information from topologically adjacent points, and combines these two methods in a sequential manner. Ju [19] adopts a self-supervised approach that learns the attention-weighted loss for each pixel point and introduces penalties for different surface areas, which can retain more detailed gradient information.
@@
Here $I_{j}$ represents the image pixel intensity in the $jth$ illumination direction. $\rho$ is the bidirectional reflectance distribution function (BRDF, an important formula in the field of optics and graphics used to describe how light is reflected from a given direction of incident light and outgoing light at a surface [20]) and $max(\bm{n}^{T}\bm{l_{j}},0)$ represents an attached shadow, and $\mu_{j}$ represents the noise of the camera and the environment. We assume that the pixel intensities of an image are normalized by the corresponding lighting intensities, so $\bm{l}$ can be viewed as a unit vector. The challenge of dealing with non-Lambertian surfaces limits the applicability of some photometric stereo methods that are suitable only for Lambertian surfaces, such as the least squares method. To address this issue, deep learning methods have emerged as a powerful tool for fitting surfaces of objects that cannot be expressed by mathematical formulas. However, even with the impressive fitting ability of deep learning methods, reconstruction results can still be fuzzy and have large angle errors due to the shape or material of some areas. As shown in Figure 2, complex scenes with intricate structures can be particularly challenging. Fortunately, this paper proposes a novel approach to address these “difficult” regions using multi-scale attention feature fusion, which is both simple and effective.
@@ 3 Methods
In the normal regression part, to prevent overfitting caused by a deep network, we added a dense-block structure [21]. This structure enhances feature propagation and reuses features from low-dimensional inputs to improve the accuracy of the network in predicting the normal image pixels while retaining the features of shallow local regions. In addition, we used batch normalization between each layer of the convolutional neural network to adjust the weights of the neurons into a standard normal distribution regularization layer. Our network architecture consists of six convolutional layers, two downsampling layers, five upsampling layers, two residual multi-scale attention feature fusion modules, one max-pooling layer, and one dense-block module. Furthermore, we utilized L2-Norm layers to normalize the surface normal vectors.
@@
$\bm{Branch_{j}^{i}}=\left\{\begin{array}[]{l}AsyConv(\bm{F_{i}}){\rm{}},{\rm{}}j=1{\rm{}}\\ AsyConv(\bm{F_{i}}\oplus\bm{Branc{h_{j-1}}}){\rm{}},{\rm{}}j=2,3,4\end{array}\right.$ (3)
@@ 4 Experimental Results
Datasets
The insufficient number of datasets is a limiting factor that affects the accuracy of photometric stereo predictions. In the experiments conducted in this study, the training datasets used were the same as those used by most all-pixel methods. The two shape datasets are the Blobby and Sculpture datasets [22, 23], which were rendered using the MERL BRDF dataset [24] containing information on 100 different materials. The two datasets contain 25,920 and 59,292 shapes, respectively, and each captured shape has 64 different lighting conditions, resulting in a total of 5.4 million images. The Sculpture dataset is more complex and contains more detailed information. During the training process, we used a ratio of 99:1 for training and validation data.
The DiLiGenT dataset [25] consists of 10 real-world objects with complex shapes and different materials, each captured in 96 lighting orientations. Ground truth information is provided for each object in the dataset. To evaluate the effectiveness of our proposed model, we used this dataset for quantitative assessment.
@@
Effectiveness of a Residual Multi-Scale Attention Feature Fusion Module
@@
“w/o RMAFF” removes the RMAFF module from the network and only concatenates shallow-deep features. “Single RMAFF” employs a single branch RMFE module to extract different levels of features from the deep image. “w/o Attention” removes the attention allocation mechanism from the module and directly connects the feature information extracted by the residual multi-scale module. “RMAFF+AFF” replaces the concat operation with the AFF structure [28] for feature stitching. Our approach uses two RMFE modules to extract different features from the shallow and deep layers and then stitches them together using the concat operation.
@@
The Validity of Normal Regression Network Structures
@@
Effect of Different Resolutions and Number of Input Images
Images of different resolutions may contain varying levels of information, with larger images often containing richer features such as geometry and texture [29]. Our proposed RMAFF module is designed to fully perceive light and dark variations within an image and extract the most appropriate image features based on those variations. However, due to GPU memory limitations, training models with larger numbers and higher resolutions of input images can significantly increase the training time.
@@
Effect of Different Training Datasets
In this study, the influence of dataset complexity on the accuracy of the proposed network model is analyzed. It is well known that the accuracy of the network model can be affected by the complexity of the dataset, as even better networks may not necessarily improve model accuracy, a phenomenon known as “Kolmogorov complexity” in machine learning [30, 31]. To investigate the effect of dataset complexity on network performance, we conducted separate experiments on three different training datasets and reconstructed the normal map of the “Buddha” object. It should be noted that everything was kept the same for the network except for the training dataset. The experimental results are shown in Figure 9. Our analysis indicates that the network’s performance in reconstructing surface normals is better when the surface complexity of the training dataset is higher (Train_Blobby vs. Train_Sculpture) and the dataset has a larger number of samples (Train_Blobby vs. Train_Blobby+Sculpture). Therefore, we conclude that the complexity of the object shape and material is essential in the photometric stereo task, enabling the network to capture more features of the object surface and achieve a more accurate normal estimation during the test phase.
@@
Benchmark Comparison of the DiLiGenT Dataset
We conducted a comprehensive comparison of the RMAFF-PSN with several commonly used methods, which included linear least squares-based methods (L2 [1]), four per-pixel methods (CNN-PS [6], LMPS [32], SPLINE-Net [33], PX-Net [7]), and four all-pixel methods (IRPS [34], PS-FCN [5], CHR-PSN [8], MF-PSN [9]). To evaluate the performance of each method, we used the DiLiGenT dataset, which included 96 input images for all methods, except for the “Bear” object, which used 76 images due to corruption in the first 20 images. The results of our experiments are presented in Table 2. Our proposed RMAFF-PSN method outperformed most of the networks, exhibiting higher accuracy than the existing deep learning methods under varying light distributions and image sizes (Figure 8b,d). Furthermore, we provide a visual comparison of the predicted normal maps for each object in the DiLiGenT dataset in Figure 10.
@@
L2 [1] 4.10 8.41 8.89 8.39 14.65 14.92 18.50 19.80 25.60 30.62 15.39
WG10 [35] 2.06 6.73 7.18 6.50 13.12 10.91 15.70 15.39 25.89 30.01 13.35
AZ08 [26] 2.71 6.53 7.23 5.96 11.03 12.54 13.93 14.17 21.48 30.50 12.61
GC10 [36] 3.21 8.22 8.53 6.62 7.90 14.85 14.22 19.07 9.55 27.84 12.00
IA14 [37] 3.34 6.74 6.64 7.11 8.77 10.47 9.71 14.19 13.05 25.95 10.60
ST14 [38] 1.74 6.12 6.51 6.12 8.78 10.60 10.09 13.63 13.93 25.44 10.30
SPLINE-Net [33] 4.51 6.49 8.29 5.28 10.89 10.36 9.62 15.50 7.44 17.93 9.63
EW20 [39] 1.58 6.30 6.67 6.38 7.26 13.69 11.42 15.49 7.80 18.74 9.53
DPSN [23] 2.02 6.54 7.05 6.31 7.86 12.68 11.28 15.51 8.01 16.86 9.41
CK18 [40] 1.50 5.74 6.24 4.97 8.64 8.86 10.00 11.44 11.33 21.90 9.06
IRPS [34] 1.47 5.44 6.09 5.79 7.76 10.36 11.47 11.03 6.32 22.59 8.83
LMPS [32] 2.40 6.11 6.54 5.23 7.48 9.89 8.61 13.68 7.98 16.18 8.41
PS-FCN [5] 2.82 6.16 7.13 7.55 7.25 7.91 8.60 13.33 7.33 15.85 8.39
Attention-PSN [27] 2.93 6.14 6.92 4.86 6.97 7.75 8.42 12.90 6.86 15.44 7.92
DR-PSN [41] 2.27 5.42 7.08 5.46 7.21 7.84 8.49 12.74 7.01 15.40 7.90
GPS-Net [26] 2.92 5.42 6.04 5.07 7.01 7.77 9.00 13.58 6.14 15.14 7.81
CHR-PSN [8] 2.26 5.97 7.04 6.35 6.76 7.15 8.32 12.52 6.05 15.32 7.77
CNN-PS [6] 2.12 4.38 5.37 4.20 6.38 8.07 7.42 12.12 7.92 13.83 7.18
PS-FCN(Norm.) [42] 2.67 4.76 6.17 7.72 7.15 7.53 7.84 10.92 6.72 12.39 7.39
PX-Net [7] 2.03 4.39 5.08 4.13 5.10 7.61 6.90 10.26 4.69 13.10 6.33
NormAttention-PSN [43] 2.93 4.65 5.96 5.48 6.42 7.12 7.49 9.93 5.99 12.28 6.83
@@
Qualitative Comparison of Other Real-World Datasets